# Screened Thin-Target Bremsstrahlung with Partially-Ionized High-Z Species

### v1

@@ V Conclusion and Outlooks
In this paper, the general formalism describing the calculation of bremsstrahlung differential cross sections was reviewed, addressing the main challenges still faced by existing theoretical models. The commonly adopted expression for the doubly differential cross section (DDCS) is based on the cross section in a pure Coulomb field, calculated within the Furry–Sommerfeld–Maue (FSM) picture and corrected for screening effects in the Born approximation. The novelty of the present work lies in a new formulation of the screening correction such that (i) the model accounts for screening in partially ionized states, and (ii) the correction is expressed in fully analytic form, allowing instantaneous evaluation and efficient exploration of large parameter spaces.
@@
The performance of the screened model of Eq. (3) in the neutral case was evaluated by comparison with experimental measurements across a range of atomic numbers, emission angles, and incident electron energies. Excellent agreement was obtained for forward emission angles and incident energies up to about 5 MeV, from low- to high-$Z$ elements. At larger emission angles, the model systematically underestimates the measured spectra, consistent with previous observations [32, 33, 21]. This discrepancy is attributed to the breakdown of the FSM approximation at larger deflection angles and becomes apparent at smaller angles as the energy increases. However, this limitation is not critical, since cross sections in these regions are several orders of magnitude smaller than in forward emission.
Addressing the identified limitations may require moving beyond the present set of approximations and solving the Dirac equation for exact electron wave functions in a screened potential. Such an extension would be computationally demanding but represents a natural direction for future work. Additionally, new and extended experimental data would be highly valuable for further benchmarking and refinement of the models.
@@ Acknowledgments
Appendix A Olsen–Maximon–Wergeland Additivity Rule
Figure 9 emphasizes the validity of the Olsen–Maximon–Wergeland additivity rule (3). The left panel depicts the logarithm of the ratio of the Sauter and screened Bethe–Heitler (multi-Yukawa) cross sections, as a function of the emitted photon energy and the emission angle. The incident electron has a kinetic energy of 4.54 MeV. The screening correction is most important in the soft-photon, low emission angle regime. The right panel, on the other hand, shows the ratio of the Roche–Ducos–Proriol and Sauter’s cross sections. The Coulomb (nuclear) effects are prominent in the forward direction and the high frequency limit.
Figure 9: Left: Region of the parameter space where the screening is prominent, that is soft-photons and forward emissions, for an incoming electron of energy $E_{0}=4.54$ MeV on a gold target. Right: Region of the parameter space where Coulomb corrections are required, hard photons and forward emissions.
Figure 10: Top: Theoretical model components for a 30 MeV electron incident on gold, with emission at $\theta_{0}=0^{\circ}$. Bottom: Same, but for $\theta_{0}=5^{\circ}$.
Appendix B Higher-Energy Results
Figure 10 shows the components of the model in Eq. (3) for a 30 MeV electron incident on a gold atom. The screened cross section is computed using a nuclear potential described by three Yukawa exponentials. The components behave consistently, with no significant discrepancies among themselves or relative to the results observed at lower energies.
The choice of 30 MeV for the incident electron is arbitrary, selected to be roughly three times higher than for the highest available experimental dataset. Combined with the observations in Section IV, these results suggest that the model in Eq. (3) can reasonably be applied at energies of this magnitude, provided the system remains in a regime where the breakdown of FSM is minimal.

### v2

@@ V Conclusion and Outlooks
In this paper, the general formalism describing the calculation of bremsstrahlung differential cross sections was reviewed, addressing the main challenges still faced by existing theoretical models. The commonly adopted expression for the doubly differential cross section (DDCS) is based on the cross section in a pure Coulomb field, calculated within the Furry–Sommerfeld–Maue (FSM) picture and corrected for screening effects in the Born approximation. The novelty of the present work lies in a new formulation of the screening correction such that (i) the model accounts for screening in partially ionized states, and (ii) the correction is expressed in fully analytic form, allowing instantaneous evaluation and efficient exploration of large parameter spaces. In addition, the Coulomb-field contribution entering this formulation was evaluated using the corrected next-to-leading-order Roche–Ducos–Proriol (RDP) expression.
@@
The performance of the screened model of Eq. (3) in the neutral case was evaluated by comparison with experimental measurements across a range of atomic numbers, emission angles, and incident electron energies. Excellent agreement was obtained for forward emission angles and incident energies up to about 5 MeV, from low- to high-$Z$ elements. At larger emission angles, the model systematically underestimates the measured spectra, consistent with previous observations [31, 32, 20]. This discrepancy is attributed to the breakdown of the FSM approximation at larger deflection angles and becomes apparent at smaller angles as the energy increases. However, this limitation is not critical, since cross sections in these regions are several orders of magnitude smaller than in forward emission. In practical situations, angular integration and the finite spatial and momentum spread of the electron beam further increase the contribution of forward emission.
Addressing the identified limitations may require moving beyond the present set of approximations and solving the Dirac equation for exact electron wave functions in the screened potential of ionized atoms. Such an extension would be computationally demanding but represents a natural direction for future work. Additionally, new and extended experimental data would be highly valuable for further benchmarking and refinement of the models.